### v1

# Unexpected consequences of Post-Quantum theories in the graph-theoretical approach to correlations

Abstract
This work explores the implications of the Exclusivity Principle (EP) in the context of quantum and post-quantum correlations. We first establish a key technical result demonstrating that given the set of correlations for a complementary experiment, the EP restricts the maximum set of correlations for the original experiment to the anti-blocking set. Based on it, we can prove our central result: if all quantum behaviors are accessible in Nature, the EP guarantees that no post-quantum behaviors can be realized. This can be seen as a generalization of the result of [Phys. Rev. A 89, 030101(R)], to a wider range of scenarios. It also provides novel insights into the structure of quantum correlations and their limitations.
I Introduction
Quantum Theory (QT) is one of the most successful scientific theories ever created by humankind; in that its predictions match with great accuracy empirical data. One paramount example is the loophole-free confirmation of the existence of quantum Bell-nonlocality [1, 2, 3, 4, 5]. Nonetheless, it remains a challenge to formulate it as a consequence of physically justifiable principles (of Nature), similar to what occurs with Einstein’s Special Relativity. Several distinct approaches have been proposed to address this problem. For example, certain approaches adopt an axiomatic framework, beginning with a more physically intuitive set of axioms from which one aims to derive the standard and abstract postulates of quantum theory (the Hilbert space structure, density matrices and measurement operators acting on it, and Born’s rule) [6, 7, 8, 9]. Another program is the one started by Popescu and Rohrlich’s seminal work in trying to answer the question: are quantum correlations singled out by the non-signaling principle [10]? The answer was found to be negative, as there exist correlations that satisfy the non-signaling condition but cannot be reproduced within the framework of quantum theory, with Popescu-Rohrlich boxes being the most prominent example [11]. Although this approach does not fully explain quantum correlations, it raises the broader question: What is the fundamental physical principle that explains quantum correlations [^1] An equivalent way of phrasing this question is: given that there exist correlations satisfying relativity, but more Bell-non-local than the quantum ones, what in Nature prohibits their existence? Or even, why is Nature not more non-local? What limits Nature’s non-locality? All things considered, this program’s idea is to study the set of quantum correlations in Bell scenarios from the outside, investigating physically motivated device-independent restrictions strong enough to rule out, at best, all post-quantum behaviors.? Several principles to address this question, named as the problem of quantum correlations, have since then been proposed: Non-Triviality of Communication Complexity [12], Macroscopically Locality [13], Local Orthogonality [14] and Information Causality [15, 16] being examples of. All those trials focus on the problem of distant correlations, with the bipartite Bell scenario used as a playground for initial discussions, which could be further expanded to multipartite scenarios. In this work we have a special interest in the so-called Exclusivity Principle (EP) [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27], which avoids the necessity of parts, dealing with the problem of quantum correlations in Kochen-Specker (KS) contextuality scenarios [28].
In 2010, Cabello, Severini, and Winter (CSW) presented an interesting connection between graph-theoretic concepts and KS-contextuality [29, 30]. This framework starts with a set of measurement events $e_{1},\ldots,e_{n}$. Two events are considered exclusive if they correspond to different outcomes of the same measurement [^2] A measurement event $e_{i}$ is to be understood as the performance of a given contextuality scenario, meaning $e_{i}=a,b,...,c|x,y,...,z$; in words, obtaining outcomes $a,b,...,c$ conditioned, respectively, on the chosen measurement context $x,y,...,z$. Two events $e=a,b,...,c|x,y,...,z$ and $e^{\prime}=a^{\prime},b^{\prime},...,c^{\prime}|x^{\prime},y^{\prime},...,z^{\prime}$ are said to be exclusive if $(x=x^{\prime}\land a\neq a^{\prime})\lor(y=y^{\prime}\land b\neq b^{\prime})\lor...\lor(z=z^{\prime}\land c\neq c^{\prime})$.. We associate each event $e_{i}$ with a vertex $i$, and we say that two vertices $i$ and $j$ are adjacent, denoted as $i\sim j$, if the corresponding measurement events $e_{i}$ and $e_{j}$ are exclusive. Based on these vertices and adjacency relations, we construct a graph $G$, which we refer to as the exclusivity graph.
For a given exclusivity graph $G$, we consider theories that assign probabilities to the measurement events corresponding to its vertices. A behavior associated with $G$ is represented by a vector $(p(e_{1}),\ldots,p(e_{n}))$, where $p(e_{i})$ represents the probability of event $e_{i}$. Since any two exclusive events correspond to different outcomes of some measurement, the behavior of a graph must also satisfy the relation $p(e_{i})+p(e_{j})\leq 1$, whenever $i\sim j$ [31].
The set of behaviors is determined by the exclusivity constraints defined by the graph $G$ and by the underlying physical theory used to describe the system. For instance, a behavior $(p(e_{1}),\ldots,p(e_{n}))$ is called deterministic noncontextual if it satisfies $p(e_{i})\in\{0,1\}$ and $p(e_{i})+p(e_{j})\leq 1$ for all pairs $i\sim j$ [32]. The set of all noncontextual (or classical) behaviors, denoted as $\mathrm{NC}(G)$, is then obtained by making probabilistic mixtures of such deterministic behaviors, a construction known as the convex hull of those deterministic noncontextual behaviors [^3] In graph-theoretical terms, this set is known as $\mathrm{STAB}(G)$ [30, 33].. Behaviors that do not belong to $\mathrm{NC}(G)$ are considered contextual.
[…]
In quantum theory, the pairwise joint measurability of a set of observables implies the joint measurability of the entire set [34]. Consequently, for quantum behaviors, if a collection of measurement events ${e_{k}}$ is pairwise exclusive, the constraint $\sum_{k}p(e_{k})\leq 1$ must hold. The Exclusivity Principle (EP) extends this property as a universal hypothesis, proposing that this constraint should be satisfied in any physical theory. In other words, the EP asserts that for any set of pairwise exclusive measurement events ${e_{k}}$—that is, events that can be associated to mutually exclusive outcomes of a single mother measurement—the sum of their probabilities must satisfy $\sum_{k}p(e_{k})\leq 1$. We define EP1-behaviors as those that adhere to the Exclusivity Principle. The set of all such behaviors is referred to as the (single-copy) EP set [^5] In graph-theoretical terms, this set is referred to as $\mathrm{QSTAB}(G)$ [30, 33]., denoted by $E_{1}(G)$.
Cabello, Severini, and Winter [29, 30] showed that given any non-contextuality (NC) expression appropriately written as a positive linear combination of probabilities of measurement events, $S_{\omega}=\sum_{i}\omega_{i}p(e_{i})$, its corresponding classical (or non-contextual), quantum, and (single-copy) exclusivity bounds match three characteristic numbers of the associated weight-exclusivity graph, a weighted-graph created from the exclusivity graph associated to the events $e_{1},\ldots,e_{n}$, and the weight-vector $\omega=(\omega_{1},\ldots,\omega_{n})$. Specifically, this relationship is expressed as
$S_{\omega}\leq^{\mathrm{NC}}\alpha(G,\mathbf{\omega})\leq^{\mathrm{Q}}\vartheta(G,\mathbf{\omega})\leq^{\mathrm{E}_{1}}\alpha^{*}(G,\mathbf{\omega}),$ (1)
in other words, $\mathrm{max}_{\mathrm{NC}(G)}S_{\mathbf{\omega}}=\alpha(G,\mathbf{\omega})$, $\mathrm{max}_{\mathrm{Q}(G)}S_{\mathbf{\omega}}=\vartheta(G,\mathbf{\omega})$ and $\mathrm{max}_{\mathrm{E}_{1}(G)}S_{\mathbf{\omega}}=\alpha^{*}(G,\mathbf{\omega})$, where $\alpha(G,\mathbf{\omega})$ is the independence number, $\vartheta(G,\mathbf{\omega})$ is its Lovász number, and $\alpha^{*}(G,\mathbf{\omega})$ is its fractional packing number, being all these three numbers well-studied objects in graph theory (and combinatorial optimization) [35, 36, 37, 33, 31].
Now, since the quantum bound of any NC inequality corresponds to the Lovász number of its associated exclusivity graph $(G,\mathbf{\omega})$, and $\vartheta(G,\mathbf{\omega})$ is obtained from the set of quantum correlations of $G$, $\mathrm{Q(G)}$, if a physical principle is able to single out the latter (or, equivalently, justify why post-quantum graph behaviors are not allowed), then such principle corresponds to an answer of why is Nature not more contextual. This line of reasoning was first proposed in 2013 by Cabello [17], bringing the problem of quantum correlations to the CSW approach.
By considering two copies of the pentagon graph (the exclusivity graph associated with the KCBS inequality, known for being the simplest NC inequality [38]), Cabello [17] was able to show that the maximum value allowed by EP to it is $\sqrt{5}$, which also equals its maximum quantum value. Shortly afterward, Yan proposed [21] a more general construction, involving the consideration of any given experiment and what is called its respective complementary experiment. This approach proved highly effective for applying the EP and enabled the extension of previous results. All results to be recalled in this paper, as well as the original ones to be shown, are derived from such underlying construction. Let us give it some special attention.
Let $\{e_{i}\}$ be a set of $n$ measurement events with exclusivity graph $G$ and $\{e_{j}^{\prime}\}$ a set of also $n$ measurement events but with exclusivity graph $\bar{G}$, where $\bar{G}$ is the complementary graph [^6] Complementary graph $\bar{G}$ is defined with the same vertex set as $G$, i.e., $V(\bar{G})=V(G)$, but its edge set is defined by relation $ij\in E(\bar{G})\iff ij\notin E(G)$. of $G$. One then defines a set of composite events $\{f_{ij}=e_{i}\land e_{j}^{\prime}\}_{ij}$ ($f_{ij}$ occurs iff $e_{i}$ and $e_{j}^{\prime}$ occurs). The exclusivity graph encoding the exclusivity structure of events $f_{ij}$ is given by the disjunctive product [^7] Given graphs $G=(V(G),E(G))$ and $H=(V(H),E(H))$, $G*H$ is defined as: $V(G*H)=V(G)\times V(H)$; edge $(i,j^{\prime})(k,l^{\prime})\in E(G*H)$ iff $ik\in E(G)$ or $j^{\prime}l^{\prime}\in E(H)$. of $G$ and $\bar{G}$, namely $G*\bar{G}$. For any $G$, $\{f_{ii}\}_{i}$ corresponds to a subset of events in $G*\bar{G}$ consisting of only pairwise exclusive events. Consequently, EP implies
$\sum_{i}p(f_{ii})=\sum_{i}p(e_{i},e_{i}^{\prime})\leq^{\mathrm{EP}}1.$ (2)
At this point, it is further assumed that the experiments $G$ and $\bar{G}$ are completely independent, allowing the joint probabilities of local events to factorize as $p(e_{i},e_{i}^{\prime})=p(e_{i})p(e_{i}^{\prime})$. Surprisingly, even under this assumption of independence, one experiment can impose constraints on the other through the EP inequality, Ineq. (2). Such a paradigm is hereafter referred to as Yan’s construction. By making assumptions about the available behavior set of (or, equivalently, theory obeyed by) the complementary experiment and applying the resulting EP inequality, several significant results were obtained.
Yan first demonstrated [21] that, given $\mathrm{Q(\bar{G})}$, EP singles out the maximum quantum value of inequality $S=\sum_{i}p_{i}$ for experiment $G$. However, it is worth mentioning that there are infinitely many other inequalities [^8] Specifically, any inequality $S_{\omega}$ with weights differing from $1$ for each measurement event., which are based on the exclusivity structure of experiment $G$, for which Yan’s result says nothing about. In 2014, Amaral, Terra, and Cabello (ATC) established a stronger result [22]: given $\mathrm{Q(\bar{G})}$, EP singles out $\mathrm{Q(G)}$. This effectively resolves the limitation of the previous finding. Since EP picks out the entire quantum behavior set, it follows that for any chosen inequality $S_{\omega}$, EP also identifies its quantum maximum. Furthermore, ATC demonstrated that when the experiment consists of self-complementary graphs [^9] A graph $G$ is self-complementary if it is isomorphic to its complement $\bar{G}$; an example is the KCBS graph., EP forbids Post-Quantum sets [^10] Where by Post-Quantum sets we mean sets strictly larger than $\mathrm{Q(G)}$.. A logical next step would be to ask whether such a result holds even without the self-complementarity assumption, i.e., to show that EP forbids sets larger than the quantum for any given experiment. In 2019, Cabello demonstrated [26, 27] through a series of intricate arguments — for instance, considering an infinite number of copies of an experiment — that EP forbids sets larger than $\mathrm{Q(G)}$ for self-complementary graphs. A second key result in the same paper provides a systematic procedure for constructing a self-complementary graph $H(G)$ from any given initial graph $G$. By combining this graph construction technique with his first result and using some additional involved results from graph theory, Cabello was ultimately able to show that EP singles out $\mathrm{Q(G)}$ for any experiment. In some sense, the 2019 work solves the quantum correlations problem within the CSW approach. However, it does so by employing complex arguments, including constructions that extend far beyond those proposed by Yan.
In this work, we use Yan’s construction together with the notion of anti-blocking duality theory [39] to show that correlation sets strictly larger than $\mathrm{Q(G)}$ are ruled out by the EP.
II Results
We begin by presenting the core technical contribution of this work in Proposition 1, where we establish that if the set of correlations for the complementary experiment is given by $X(\bar{G})$, then the EP constrains the maximum set of correlations for the original experiment to the anti-blocking set of $X(\bar{G})$. A direct consequence of this proposition, combined with a foundational graph-theoretical result [33], leads to our first result: assuming the set of correlations for the complementary experiment is the $\mathrm{E_{1}(G)}$, the EP restricts the set of correlations for the experiment to the classical set (Corollary 1). Lastly, we present the central result of this paper: quantum theory and EP exclude post-quantum theories. This can be seen as a generalization of ATC’s second result for any exclusivity graph. This generalization addresses the previously identified logical gap in the literature and provides valuable insights into the findings of the 2019 work.
Let us start with a remark.
Remark 1.
Any behavior $p^{\prime}$ achievable in the complementary experiment $\bar{G}$, imposes a constraint on the independent experiment $G$.
Indeed, under Yan’s construction, EP inequality (2) becomes
$\sum_{i}p(e_{i})p(e_{i}^{\prime})\leq 1.$ (3)
The above inequality can be seen as a half-space separation over the set of behaviors for the experiment $G$, where the probabilities $p(e_{i}^{\prime})$ are seen as the coefficients of this inequality. Moreover, we can progressively consider broader sets of achievable behaviors for the complementary experiment, and subsequently investigate the behavior sets of the experiment as defined by all resulting EP constraints of Ineq. (3). For the remainder of the paper, we will refer to these graph correlation sets as theories for experiments.
Proposition 1.
Given that the theory describing the complementary experiment is $X(\bar{G})$, the largest theory allowed by EP for the experiment is $Y(G)=\mathrm{abl}X(\bar{G})$.
Proof.
On the one hand, $Y(G)$ is defined by all EP inequalities in form of Ineq. (3), which arise when considering all behaviors of the complementary experiment $\bar{G}$; this can be expressed mathematically as follows:
$Y(G)=\{p\geq 0\>|\>\sum_{i}p_{i}p_{i}^{\prime}\leq 1\>\forall\>p^{\prime}\in X(\bar{G})\}.$ (4)
On the other hand, for any set of non-negative vectors $A$, its anti-blocking associated set, denoted $\mathrm{abl}A$, is given by
$\mathrm{abl}A=\{b\geq 0\>|\>\sum_{i}a_{i}b_{i}\leq 1\>\forall\>a\in A\}.$ (5)
Straightforward comparison between Eq. (4) and Eq. (5) reveals that, under Yan’s construction, EP constraint Ineq. (3) matches exactly the defining property of an anti-blocking set. Therefore, $\mathrm{Y(G)}=\mathrm{abl}X(\bar{G})$. ∎
Although Proposition 1 holds for all theories $X(\bar{G})$, we will focus on the most relevant cases. These range from the smallest possible set to describe $\bar{G}$, corresponding to classical theory, $\mathrm{NC}(\bar{G})$, to the largest possible set consistent with EP, corresponding to a theory that only demands single-copy exclusivities, denoted $\mathrm{E_{1}}(\bar{G})$.
In the following, we investigate the restrictive power of EP in selecting theories for a given experiment, by assuming specific theories for the complementary experiment, including the limiting cases of classical and $E_{1}$ theories, as well as QT.
Corollary 1.
Given $\mathrm{E_{1}(\bar{G})}$, EP singles out $\mathrm{NC(G)}$. Conversely, given $\mathrm{NC(\bar{G})}$, EP singles out $\mathrm{E_{1}(G)}$.
Proof.
By Proposition 1, the largest theory allowed by EP to the experiment is $Y(G)=\mathrm{abl\>}\mathrm{E_{1}(\bar{G})}$. However, as shown in Ref. [33], for any experiment, $E_{1}$ and classical theories are connected via the anti-blocking operation, specifically $\mathrm{E_{1}(\bar{G})}=\mathrm{abl\>}\mathrm{NC(G)}$. Consequently, it follows that $\mathrm{abl\>E_{1}(\bar{G})}=\mathrm{abl\>abl\>NC(G)}$. Since for any convex corner [^11] A set of vectors $V$ is a convex corner if it is nonempty, closed, convex, nonnegative and $0\leq\vec{u}\leq\vec{v}$ and $\vec{v}\in V\Rightarrow\vec{u}\in V$; see section $30$ of [33]. $D$, we know that $\mathrm{abl\>abl\>}D=D$ and both $\mathrm{E_{1}(\bar{G})}$ and $\mathrm{NC(G)}$ are convex corners [33], it results that $Y(G)=\mathrm{abl\>E_{1}(\bar{G})}=\mathrm{NC(G)}$. As for the second part, the distinction between the label of an experiment and its complement is entirely arbitrary. Therefore, by exchanging $G$ with $\bar{G}$, we obtain $\mathrm{E_{1}(G)}=\mathrm{abl\>NC(\bar{G})}$. ∎
The first part of Corollary 1 states that if EP holds and there exists a complementary experiment $\bar{G}$ where it is provable that all EP1-behaviors are accessible [^12] The case of interest being an imperfect graph; for perfect graphs $\mathrm{NC(\bar{G})=Q(\bar{G})=E_{1}(\bar{G})}$., then to the experiment $G$, only noncontextual behaviors are allowed. As for the second part, if EP is the only restriction and classical theory is accessible to $\bar{G}$, then to the experiment $G$, all EP1-behaviors are allowed. Notably, the larger the theory assumed for the complementary experiment, the stronger EP becomes in excluding behaviors from the experiment. Conversely, the smaller the theory assumed for the complementary experiment, the weaker EP becomes in excluding behaviors from the experiment. This mechanism is precisely what underpins the proof of our main result. Before showing it, we remark that Proposition 1 recovers result $1$ due to ATC in Ref. [22].
Remark 2.
Given $\mathrm{Q(\bar{G})}$, EP singles out $\mathrm{Q(G)}$.
Proof.
By Proposition 1, the largest theory allowed by EP to the experiment is $Y(G)=\mathrm{abl\>}\mathrm{Q(\bar{G})}$. As shown in Ref. [33], for any experiment the equality $\mathrm{Q(\bar{G})}=\mathrm{abl\>Q(G)}$ holds. Consequently, $\mathrm{abl\>Q(\bar{G})}=\mathrm{abl\>abl\>Q(G)}$, and since both $\mathrm{Q(G)}$ and $\mathrm{Q(\bar{G})}$ are convex corners, it follows that $Y(G)=\mathrm{abl\>Q(\bar{G})}=\mathrm{Q(G)}$. ∎
Moreover, we can see Proposition 1 as a generalization of Result $1$ of Ref. [22] framed in a theory-independent manner, as it determines the largest set allowed by EP for the experiment, $Y(G)=\mathrm{abl}\>X(\bar{G})$, for any assumed theory describing the complementary experiment, $X(\bar{G})$, ranging from the noncontextual to the $E_{1}$ theory.
An intriguing fact arises from the result $\mathrm{Q(\bar{G})}=\mathrm{abl\>Q(G)}$ used in the previous proof. In essence, the theory assumed for the complementary experiment, quantum, is precisely the same theory singled out by EP. This stands in contrast to the case where classical or $E_{1}$ theory were assumed (recall that if classical theory were the sole framework for describing physical phenomena in a complementary experiment, EP, by singling out $E_{1}$ theory for the experiment, asserts that the latter is governed by a completely different set of rules— or, in other words, distinct physical laws). That a theory is consistently utilized to describe each and every experiment is certainly a desirable property, for easy-to-understand reasons. In this context, the following question arises: are there other theories $U$ that satisfy the self-duality equation $\mathrm{U(\bar{G})}=\mathrm{abl\>U(G)}$? Should this be the case, Proposition 1 would yield the same conclusions as those drawn from the quantum scenario. Furthermore, as far as EP can ascertain, both theories would be equally valid in describing empirical data. We claim that if the uniqueness of QT as the solution to the self-duality equation is demonstrated, then it is shown EP explains QT within the CSW framework. A partial answer to the above question is now discussed.
It seems adequate to argue that for such hypothetical theory, $U$, to be of physical interest and potentially lead to new physics—for instance, by allowing the possibility of violating quantum bounds of NC inequalities—it must extend beyond QT while still encompassing all of its predictions. In the following, we argue that if such a theory exists, it would imply that certain quantum behaviors are prohibited in an independent experiment.
Main Result.
If every quantum behavior is experimentally accessible, then EP excludes post-quantum behaviors. Equivalently, quantum theory and EP exclude post-quantum theories.
Proof.
We will prove the contrapositive. That is, if the set of behaviors (or the theory) associated with the experiment $\bar{G}$ is post-quantum, then this implies that the EP restricts the set of behaviors (or the theory) for the experiment $G$, such that it becomes strictly contained within the set of quantum behaviors. In other words, under these conditions, certain quantum behaviors would become experimentally inaccessible. Let then be $U(\bar{G})\equiv\mathrm{Q(\bar{G})}\cup W(\bar{G})$ the set behaviors for the experiment $\bar{G}$, where each $\omega^{\prime}\in W(\bar{G})$ corresponds to a post-quantum behaviour; i.e., $\omega^{\prime}\in\mathrm{E_{1}(\bar{G})}\setminus\mathrm{Q(\bar{G})}$. We assume that $U(\bar{G})$ is a convex corner that satisfies the self-duality equation $U(\bar{G})=\mathrm{abl\>}U(G)$ (though the main result still holds even in the absence of latter assumption; no further assumptions are made regarding $W(\bar{G})$, other than its non-emptiness). Thus, our post-quantum theory is represented by $U(\bar{G})$. From Proposition 1, the theory singled out by EP to describe the experiment is given by $Y(G)=\mathrm{abl\>}U(\bar{G})=\mathrm{abl\>}\{Q(\bar{G})\cup W(\bar{G})\}$. More explicitly,
$Y(G)=\Big{\{}p\geq 0|\sum_{i}p_{i}p_{i}^{\prime}\leq 1\\ \>\forall\>p^{\prime}\in U(\bar{G})=\mathrm{Q(\bar{G})\cup W(\bar{G})}\Big{\}}.$ (6)
Therefore,
$Y(G)=\Big{\{}p\geq 0\>|\>\big{[}\ \sum_{i}p_{i}q_{i}^{\prime}\leq 1\>\forall\>q^{\prime}\in\mathrm{Q(\bar{G})}\big{]}\ \\ \land\big{[}\ \sum_{i}p_{i}w_{i}^{\prime}\leq 1\>\forall\>w^{\prime}\in W(\bar{G})\big{]}\ \Big{\}},$ (7)
and finally,
$Y(G)=\mathrm{Q(G)}\cap\Big{\{}p\geq 0\>|\>\sum_{i}p_{i}w_{i}^{\prime}\leq 1\>\forall\>w^{\prime}\in W(\bar{G})\Big{\}}.$ (8)
From Eq. (8), it follows immediately that $Y(G)\subseteq\mathrm{Q(G)}$. Using the self-duality property of QT, we have
$\mathrm{Q(\bar{G})}=\Big{\{}p^{\prime}(\bar{G})\geq 0|\sum_{i}p_{i}^{\prime}p_{i}\leq 1\>\forall\>p\in\mathrm{Q(G)}\Big{\}};$ (9)
for any $p^{\prime}(\bar{G})\in\mathrm{E_{1}(\bar{G})\setminus Q(\bar{G})}$, there exists $p\in\mathrm{Q(G)}$ such that
$\sum_{i}p^{\prime}_{i}p_{i}>1.$ (10)
Let $p^{\prime}(\bar{G})=w^{\prime}\in W(\bar{G})$; then there exists $p\in\mathrm{Q(G)}$ such that
$\sum_{i}p_{i}w_{i}^{\prime}>1.$ (11)
Any such $p(G)$ is excluded from the definition of $Y(G)$ in Eq. (8) and, as a matter of fact, $Y(G)=\mathrm{abl\>U(\bar{G})}\subset\mathrm{Q(G)}$. It is even possible to show that all excluded behaviors are post-classical (or, if you will, genuinely quantum), that is, $p(G)\in\mathrm{Q(G)\setminus NC(G)}$. Indeed, from the duality between classical and $E_{1}$ theory used in corollary 1, we have
$\mathrm{NC(G)}=\mathrm{abl\>E_{1}(\bar{G})}=\Big{\{}p\geq 0|\sum_{i}p_{i}p_{i}^{\prime}\leq 1\\ \>\forall\>p^{\prime}\in\mathrm{E_{1}(\bar{G})}\Big{\}}.$ (12)
As a consequence, for all $w^{\prime}\in W(\bar{G})$, any $p(G)\in\mathrm{NC(G)}$ satisfies $\sum_{i}p_{i}w_{i}\leq 1$.
Any Post-Quantum theory assumed for the complementary experiment leads to a theory strictly smaller than Quantum for the experiment, singled out by EP. ∎
Let’s now explore the implications of our main result. Under the assumptions that EP holds and Post-Quantum theories are allowed for at least one complementary experiment, our findings ensure that some genuinely quantum behaviors become prohibited. However, quantum theory has been extensively tested in various experimental contexts, and its predictions have been found to agree with observation across an extraordinarily wide range of scenarios [40, 41, 42, 43, 1, 2, 3]. While certain quantum behaviors may be technically challenging to realize [44, 45, 46, 47], there is no fundamental reason to believe that any of them are impossible to achieve in Nature. Therefore, it is plausible to argue that one of the two assumptions must be false.
Given that EP is our chosen physical principle, we will first assume that EP is indeed respected in Nature—a view supported by experimental evidence [19]. Consequently, we conclude that EP imposes a fundamental limitation, ruling out the existence of theories that extend beyond Quantum Theory. This conclusion generalizes ATC’s second result in [22] to arbitrary experiments, beyond the specific case of self-complementary graphs.
Continuing with the assumption that no quantum behavior is fundamentally impossible to realize in Nature, an alternative perspective for those who do not view QT as the ultimate theory of Nature is that the EP does not hold universally. In other words, if a Post-Quantum behavior were ever observed in an experiment, our result would ensure a breakdown of the Exclusivity Principle.
Finally, the question of whether there exist theories $U(\bar{G})$ that satisfy the self-duality equation can now be interpreted as an enhancement of our main result. Indeed, if it can be proven that QT is the only theory with such a property, then any alternative theory would fail to be universal, as some of its predictions would be prohibited in independent experiments. This would demonstrate that EP uniquely singles out QT, without presupposing that all quantum behaviors are necessarily realizable in Nature. Moreover, our result advances the resolution of this problem by showing that a theory satisfying the self-duality equation can neither strictly contain quantum theory nor be strictly contained by it.
III Conclusion
In this work, we extended previous results on the quantum correlations problem within the CSW framework. Specifically, we examined a type of well-known composite correlation experiment, Yan’s construction. We identified that the largest theory allowed by the Exclusivity Principle (EP) to describe an experiment, given a theory for the complementary experiment, aligns with the mathematical structure of anti-blocking duality. This insight enabled us to determine, for any assumed theory of any complementary experiment, the corresponding theory selected by EP for the original experiment (Proposition 1). We then explored the implications of this result for classical and $E_{1}$ theories (Corollary 1), and more importantly, for Quantum Theory (Remark 2), where we recovered a previously established result from the literature.
Building on Proposition 1 we derived our main result, where we demonstrated that Quantum Theory and EP forbid Post-Quantum theories to produce results distinct from the quantum ones. In other words, by assuming that EP holds, we derive a fundamental limitation: EP excludes the possibility of theories extending beyond Quantum Theory, generalizing previous results to arbitrary experiments. Alternatively, if Post-Quantum behaviors are observed and the corresponding Sub-Quantum counterpart cannot be certified, this would indicate that EP is not a valid universal principle.
Finally, we propose that another way to show that EP uniquely selects QT for any experiment is by proving that QT is the unique solution to the self-duality equation $U(\bar{G})=\mathrm{abl\>U(G)}$. Our work also serves as a bridge between the findings of ATC’s 2014 work [22] (along with earlier works that follow the same logical structure), by generalizing some of its results, and Cabello’s 2019 work [26], by reaching a similar conclusion for the quantum correlations problem. While both approaches highlight the role of EP, our approach reaches this result through Yan’s more streamlined construction, offering a simpler framework.
Acknowledgements.
We thank Bárbara Amaral, Adán Cabello and Rafael Rabelo for enlightening discussions. This study was financed in part by the Brazilian National Council for Scientific and Technological Development (CNPq) and in part by the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior - Brasil (CAPES) - Finance Code 001. This work was also supported by CNPq under grant no. 311314/2023-6, and it is part of the Brazilian National Institute for Science and Technology in Quantum Information (INCT-QI).